OpenStudy (anonymous):

Find the quotient. 16 divided by 1/4 Choose exactly two answers that are correct.

OpenStudy (fibonaccichick666):

SO do you know the chant for dividing fractions?

OpenStudy (anonymous):

a= 64/1 b=64 c=4/1 d=4

OpenStudy (fibonaccichick666):

the chant, which you should memorize is "Flip the second fraction and.. MULTIPLY!

OpenStudy (fibonaccichick666):

so the first think you want to do is turn the 16 into a fraction, can you do that?

OpenStudy (anonymous):

ys

OpenStudy (fibonaccichick666):

please do then

OpenStudy (anonymous):

1/16

OpenStudy (fibonaccichick666):

that is 1 sixteenth, I don't think that is the same as 16, try again?

OpenStudy (anonymous):

never mine im closing

OpenStudy (fibonaccichick666):

why?